OpenStudy (anonymous):

3. Could I find the two missing angle measures if I know some of the side lengths of a right triangle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

um

OpenStudy (anonymous):

use Pythagoras theorem \[b^{2}+c^2 = a^2\]

OpenStudy (anonymous):

ohh yea ty

OpenStudy (anonymous):

also you can use 'sohcahtoa'

OpenStudy (anonymous):

you you know what i mean?

OpenStudy (anonymous):

yea sin tan and cos

OpenStudy (anonymous):

Pythagoras is useful to get lengths of the sides and sin cos and tan can be used to get the angle's or side length also.

OpenStudy (anonymous):

sin cos and tan there are three variables the angle and two sides as long as you have two values you can solve for the other.

OpenStudy (anonymous):

and you also know that if its a right triangle the sum of angle adds to 180 and one angle is always 90.

OpenStudy (anonymous):

ok kool ty

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

say the oposite and the base are 4.25 whats the hypotenuse length like how do i work it out

OpenStudy (anonymous):

well 4.25 = a 4.25=b and h=? \[a^2+b^2=h^2 \]

OpenStudy (anonymous):

solve for h

OpenStudy (anonymous):

is one way.

OpenStudy (anonymous):

um 36.125

OpenStudy (anonymous):

i feel like i did it wrong

OpenStudy (anonymous):

so you have \[36.125=h^2\]

OpenStudy (anonymous):

yea it just doesnt seem right

OpenStudy (anonymous):

its right

OpenStudy (anonymous):

you need to get rid of the square

OpenStudy (anonymous):

so remember you can do what every you want to on the right of the equation as long as you do it to the left.

OpenStudy (anonymous):

so it a trianlge automaticly equals 180 and i have 90 and 36 which equals 126 then the other angle is 54

OpenStudy (anonymous):